The owner of At the Beach is forecasting this month's (October's) demand for a new tanning booth based on the historical data given below.

Month

Number of Visits

April

100

May

140

June

110

July

150

August

120

September

160

a) What is this month’s forecast using the naïve approach?

b) What is this month’s forecast using a three-month simple moving average?

c) What is this month's forecast using a four-month weighted moving average with weights of .4, .3, .2, and .1?

d) What is this month's forecast using exponential smoothing with alpha = .2, if August's forecast was 145?
Answer #1
1 B C D E F
2 Month Number of Visits Forecast using Naïve approach Forecast using three month simple moving average Forecast using four month weighted moving average
3 April 100
4 May 140 100
5 June 110 140
6 July 150 110 116.67
7 August 120 150 133.33 131
8 September 160 120 126.67 129
9 October 160 143.33 141
10
11 Month Number of Visits Forecast using Exponential Smoothing
12 April 100
13 May 140
14 June 110
15 July 150
16 August 120 145
17 September 160 140
18 October 144

Formula:

1 B C D E F
2 Month Number of Visits Forecast using Naïve approach Forecast using three month simple moving average Forecast using four month weighted moving average
3 April 100
4 May 140 =C3
5 June 110 =C4
6 July 150 =C5 =AVERAGE(C3:C5)
7 August 120 =C6 =AVERAGE(C4:C6) =(0.4*C6)+(0.3*C5)+(0.2*C4)+(0.1*C3)
8 September 160 =C7 =AVERAGE(C5:C7) =(0.4*C7)+(0.3*C6)+(0.2*C5)+(0.1*C4)
9 October =C8 =AVERAGE(C6:C8) =(0.4*C8)+(0.3*C7)+(0.2*C6)+(0.1*C5)
10
11 Month Number of Visits Forecast using Exponential Smoothing
12 April 100
13 May 140
14 June 110
15 July 150
16 August 120 145
17 September 160 =(0.2*C16)+(0.8*D16)
18 October =(0.2*C17)+(0.8*D17)
Answer #2
1 B C D E F
2 Month Number of Visits Forecast using Naïve approach Forecast using three month simple moving average Forecast using four month weighted moving average
3 April 100
4 May 140 100
5 June 110 140
6 July 150 110 116.67
7 August 120 150 133.33 131
8 September 160 120 126.67 129
9 October 160 143.33 141
10
11 Month Number of Visits Forecast using Exponential Smoothing
12 April 100
13 May 140
14 June 110
15 July 150
16 August 120 145
17 September 160 140
18 October 144

Formula:

1 B C D E F
2 Month Number of Visits Forecast using Naïve approach Forecast using three month simple moving average Forecast using four month weighted moving average
3 April 100
4 May 140 =C3
5 June 110 =C4
6 July 150 =C5 =AVERAGE(C3:C5)
7 August 120 =C6 =AVERAGE(C4:C6) =(0.4*C6)+(0.3*C5)+(0.2*C4)+(0.1*C3)
8 September 160 =C7 =AVERAGE(C5:C7) =(0.4*C7)+(0.3*C6)+(0.2*C5)+(0.1*C4)
9 October =C8 =AVERAGE(C6:C8) =(0.4*C8)+(0.3*C7)+(0.2*C6)+(0.1*C5)
10
11 Month Number of Visits Forecast using Exponential Smoothing
12 April 100
13 May 140
14 June 110
15 July 150
16 August 120 145
17 September 160 =(0.2*C16)+(0.8*D16)
18 October =(0.2*C17)+(0.8*D17)